# Community detection in heterogeneous signed networks

### v1

@@ Abstract
Network data has attracted growing interest across scientific domains, prompting the development of various network models. Existing network analysis methods mainly focus on unsigned networks, whereas signed networks, consisting of both positive and negative edges, have been frequently encountered in practice but much less investigated. In this paper, we formally define strong and weak balance in signed networks, and propose a signed block $\beta$-model, which is capable of modeling strong- and weak-balanced signed networks simultaneously. We establish the identifiability of the proposed model by leveraging properties of bipartite graphs, and develop an efficient alternating updating algorithm to optimize the resulting log-likelihood function. More importantly, we establish the asymptotic consistencies of the proposed model in terms of both probability estimation and community detection. Its advantages are also demonstrated through extensive numerical experiments and the application to a real-world international relationship network.
@@ 1 Introduction
Network data has attracted increasing attention in recent years, with applications in diverse scientific domains, ranging from social science [hunter2008goodness], computer science [heard2014filtering] to biomedicine [shojaie2009analysis]. Analysis of network data can reveal latent structures underlying the interactions among entities of interest, such as community detection [girvan2002community], link prediction [martinez2016survey], homophily [hoff2007modeling] and so on.
Most existing methods for network analysis focus on unsigned networks, whereas signed networks are also prevalent in real-world applications. Specifically, signed networks have been extensively employed to represent relationships among objects [harary1953notion], which associate polarity information with each edge and thus distinguish friends from enemies, or allies from rivals. These signed edges give rise to unique structural characteristics grounded in social psychology theories, particularly the balance theory [heider1946attitudes; cartwright1956structural], stating that certain triadic configurations involving signed edges lead to stable or unstable social structures. Figure 1 illustrates four possible triads, among which triads A and C represent strong-balanced structures. These triads align with intuitive social principles, such as “a friend of my friend is likely my friend” and “an enemy of my friend is likely my enemy”. Weak balance [davis1967clustering] extends this concept by additionally permitting triad D, which accommodates the occurrence of three mutually hostile nodes.
@@
Community detection has been widely studied for unsigned networks [holland1983stochastic; girvan2002community; newman2006modularity; lei2015consistency; zhang2022directed], but much less so for signed networks. Most existing community detection methods for signed networks [doreian1996partitioning; kunegis2010spectral; jiang2015stochastic; cucuringu2021regularized] are algorithm based and lack formal theoretical analysis, not to mention that the prevalent balanced structures in signed networks are often ignored by the aforementioned community detection methods. Note that the balance theory actually posits that a strong-balanced (weak-balanced) network can be partitioned into two (multiple) sub-networks such that positive edges connect nodes within the same sub-network, whereas nodes from different sub-networks are connected by negative edges [easley2010networks]. More recently, tang2024population introduces a formal definition of strong-balanced signed networks in a probabilistic sense and proposes an efficient method to detect the two communities, yet the method is not readily applicable to the weak-balanced signed networks with multiple communities.
@@
In this paper, we first extend the definition of strong-balanced signed networks in tang2024population to accommodate weak-balanced signed networks. We also provide a set of equivalence conditions for both strong- and weak-balanced signed networks, from both local and global perspectives. Built upon the global perspective, we then propose a signed block $\beta$-model to explicitly incorporate the strong or weak balance structures in heterogeneous signed networks. Particularly, it introduces two sets of parameters to capture each node’s distinct connection tendencies within its own community and across different communities. The identifiability of the signed block $\beta$-model is established by exploiting properties of the bipartite graph, and an efficient alternating updating algorithm is developed to optimize the resulting negative log-likelihood function for parameter estimation. More importantly, the asymptotic consistencies of the block $\beta$-model are established in terms of both probability estimation and community detection. Its advantages are also supported in numerical experiments on various synthetic signed networks and a real-world international relationship network.
@@
The main contribution of the proposed signed block $\beta$-model is that it provides a novel framework to statistically model strong and weak balance structures in signed networks. It also generalizes the traditional $\beta$-model [chatterjee2011random; 10.1214/12-AOS1078; chen2021analysis] by incorporating community structures while preserving its capacity to capture node heterogeneity. For each node, two sets of parameters are introduced to model its tendency to positively or negatively connect with other nodes within the same community or across different communities. Such a treatment not only incorporates community structure in each node’s connection tendency, but also admits more substantial node heterogeneity than existing models [chen2021analysis; shao20212; cucuringu2021regularized]. Moreover, a novel set of identifiability conditions for the proposed signed block $\beta$-model is established by leveraging techniques from the graph theory. The theoretical techniques developed here can be applicable to identifiability analyses of other network models involving community structures, and thus are of independent interest.
@@
The remainder of this paper is organized as follows. Section 2 formally gives the definition of strong and weak balance in signed networks. Section 3 presents the signed block $\beta$-model, establishes its identifiability conditions, and develops an efficient alternative updating scheme to tackle the resulting optimization task. Section 4 provides a rigorous theoretical analysis of the proposed signed block $\beta$-model and quantifies its asymptotic behaviors in both probability estimation and community detection. Section 5 examines the numeric performance of the signed block $\beta$-model via extensive experiments on both synthetic and real-world signed networks. Finally, Section 6 concludes the paper with a brief summary. Technical proofs of the main theorems are provided in the Appendix, whereas auxiliary lemmas are contained in the Supplementary Material.
@@
We introduce some notations to be used throughout the paper. For any integer $n$, let $n]=\{1,\dots,n\}$. For any vector $\bm{v}=(v_{1},\dots,v_{n})^{\top}\in\mathcal{R}^{n}$, $\|\bm{v}\|_{2}^{2}=\sum_{i\in[n]}v_{i}^{2}$, and $\|\bm{v}\|_{\infty}=\max_{i\in[n]}|v_{i}|$. We use $\bm{1}_{n}$ to denote the $n$-dimensional column vector with all elements being 1 and $\mathbb{I}(\cdot)$ to denote the indicator function. For any matrix $\bm{A}=(a_{ij})\in\mathcal{R}^{n\times m}$, let $\bm{A}_{i}$ be its $i$-th row, and $\|\bm{A}\|_{F}^{2}=\sum_{i\in[n],j\in[m]}a_{ij}^{2}$ be its Frobenius norm. For any two matrices $\bm{A}$ and $\bm{B}$, $\bm{A}\odot\bm{B}$ is the Hadamard product with $(\bm{A}\odot\bm{B})_{ij}=a_{ij}b_{ij}$. For any two sequences $\{a_{n}\}_{n\geq 1}$ and $\{b_{n}\}_{n\geq 1}$, we denote $a_{n}\prec b_{n}$ if $a_{n}=o(b_{n})$, $a_{n}\preceq b_{n}$ if $a_{n}=O(b_{n})$ and $a_{n}\simeq b_{n}$ if $a_{n}\preceq b_{n}$ and $b_{n}\preceq a_{n}$. For any set $\mathcal{X}$, $|\mathcal{X}|$ denotes its cardinality and $\mathcal{X}^{c}$ denotes its complement. For any two sets $\mathcal{X}$ and $\mathcal{Y}$, $\mathcal{X}\setminus\mathcal{Y}=\{x\mid x\in\mathcal{X}\text{ and }x\notin\mathcal{Y}\}$.
2 Signed network and balance theory
@@
Definition 1 (Strong balance)
@@
Definition 1 assures that a strong-balanced signed network does not contain a triad achieving the highest probability at triad B or D as shown in Figure 1. An equivalent definition has also been introduced in tang2024population, which quantifies the probabilistic strong balance for signed networks. Moreover, Definition 1 can be extended to define weak-balanced signed networks.
Definition 2 (Weak balance)
@@
Propositions 1 and 2 characterize the strong and weak balance structures in signed networks from both local and global perspectives. In fact, the local perspective (i) in Proposition 1 is essentially equivalent to $\mathbb{E}(a_{ij}a_{jk}a_{ki}\big||a_{ij}a_{jk}a_{ki}|=1)>0$ for any $(i,j,k)$, which is exactly the definition of the strong-balanced signed network in tang2024population, whereas (i) in Proposition 2 can be regarded as its analogy for the weak-balanced signed network. On the other hand, (ii) in Proposition 1 and (ii) in Proposition 2 give the global perspectives on strong and weak balance structures, which greatly facilitate the modeling of community structures in signed networks.
@@ 3 Signed block $\beta$-model
We propose a signed block $\beta$-model, denoted as SBBM, to leverage the global perspectives of the balance structures in signed networks. Particularly, let $\psi_{i}\in[K]$ denote the community membership of node $i$, then
@@
It can be readily verified that nodes $i$ and $j$ belong to the same community if and only if $\mathbb{P}(a_{ij}=1)>\mathbb{P}(a_{ij}=-1)$, and different communities if and only if $\mathbb{P}(a_{ij}=1)<\mathbb{P}(a_{ij}=-1)$. It follows from Propositions 1 and 2 that SBBM with constraints (2) satisfies strong balance when $K=2$ and weak balance when $K\geq 3$.
@@ 3.1 Model identifiability
Let $\bm{Z}=(\bm{z}_{1},\dots,\bm{z}_{n})^{\top}$ be the community membership matrix, where $\bm{z}_{i}\in\{0,1\}^{K}$ with $z_{ik}=1$ if $\psi_{i}=k$. Define $\bm{\beta}^{+}=(\beta^{+}_{i})_{i=1}^{n}$, $\bm{\gamma}^{+}=(\gamma_{i}^{+})_{i=1}^{n}$, $\bm{\beta}^{-}=(\beta^{-}_{i})_{i=1}^{n}$, $\bm{\gamma^{-}}=(\gamma_{i}^{-})_{i=1}^{n}$, $\bm{\eta}^{+}=\bm{\beta}^{+}-\bm{\gamma}^{+}$ and $\bm{\eta}^{-}=\bm{\beta}^{-}-\bm{\gamma}^{-}$. Then, the parameter space is
$\mathcal{P}=\big\{(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z})\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}$ $\times\mathbb{R}^{n}\times\{0,1\}^{n\times n}:$
@@ 3.2 Computing algorithm
The negative log-likelihood function of SBBM is
$L(\bm{\Theta};{\bm{A}})=-\frac{2}{n(n+1)}\sum_{1\leq i\leq j\leq n}\Big(a^{+}_{ij}\theta^{+}_{ij}+a^{-}_{ij}\theta^{-}_{ij}-\log\big(1+e^{\theta^{+}_{ij}}+e^{\theta^{-}_{ij}}\big)\Big),$ (3)
where $\bm{\Theta}=(\bm{\Theta}^{+},\bm{\Theta}^{-})$, $a_{ij}^{+}=\mathbb{I}(a_{ij}=1)$, $a_{ij}^{-}=\mathbb{I}(a_{ij}=-1)$ for each $1\leq i\leq j\leq n$. Define
$\mathcal{P}_{0}=\big\{(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z})\in\mathcal{P}:\$$\eta_{i}^{+}+\eta_{j}^{+}=0\ \text{or}\ \eta_{i}^{-}+\eta_{j}^{-}=0$
$\text{for any}\ i\in S_{1},j\in S_{2}\ \text{and}\ i\in S_{1}^{c},j\in S_{2}^{c}\big\},$
where $S_{1}\subseteq C_{1}$ and $S_{2}\subseteq C_{2}$, and
[…]
Then, $\widehat{\bm{\gamma}}^{+}$, $\widehat{\bm{\eta}}^{+}$, $\widehat{\bm{\gamma}}^{-}$, $\widehat{\bm{\eta}}^{-}$ and $\widehat{\bm{Z}}$ can be obtained by minimizing $L(\bm{\Theta};{\bm{A}})$ over $(\mathcal{P}\setminus\mathcal{P}_{0})\cap\{\eta_{1}^{+}=\eta_{1}^{-}=0\}$ when $K=2$ and over $\mathcal{P}\setminus\mathcal{P}_{1}$ when $K\geq 3$. It is clear that both $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ are zero sets under the Lebesgue measure. Therefore, we get $\widehat{\bm{\gamma}}^{+}$, $\widehat{\bm{\eta}}^{+}$, $\widehat{\bm{\gamma}}^{-}$, $\widehat{\bm{\eta}}^{-}$ and $\widehat{\bm{Z}}$ by minimizing $L(\bm{\Theta};{\bm{A}})$ over $\mathcal{P}\cap\{\eta_{1}^{+}=\eta_{1}^{-}=0\}$ when $K=2$, and minimizing $L(\bm{\Theta};{\bm{A}})$ over $\mathcal{P}$ directly when $K\geq 3$ [latouche2011overlapping; zhang2022identifiability].
We develop an alternative updating scheme to optimize $L(\bm{\Theta};{\bm{A}})$. We first rewrite $L(\bm{\Theta};\bm{A})$ as $L(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z})$ and initialize $\bm{\gamma}^{+(0)}=\bm{\eta}^{+(0)}=\bm{\gamma}^{-(0)}=\bm{\eta}^{-(0)}=\bm{0}_{n}$ and $\bm{Z}^{(0)}$ by applying spectral clustering algorithm to $\bm{A}$. Given $(\bm{\gamma}^{+(t-1)},\bm{\eta}^{+(t-1)},\bm{\gamma}^{-(t-1)},\bm{\eta}^{-(t-1)},\bm{Z}^{(t-1)})$ at the tt‑th step for $t\geq 1$, note that, with $\bm{Z}^{(t-1)}$ fixed, minimizing $L(\bm{\Theta};\bm{A})$ subject to (2) with respect to $(\bm{\gamma}^{+(t-1)},\bm{\eta}^{+(t-1)},\bm{\gamma}^{-(t-1)},\bm{\eta}^{-(t-1)})$ is a convex optimization problem. We therefore adopt the Spectral Projected Gradient method [birgin2000nonmonotone], which combines a Barzilai–Borwein spectral step, Euclidean projection onto the feasible set, and an Armijo back-tracking line search, thereby handling curvature adaptation and step-size control automatically. Upon convergence, we obtain the updated parameters $(\bm{\gamma}^{+(t)},\bm{\eta}^{+(t)},\bm{\gamma}^{-(t)},\bm{\eta}^{-(t)})$. Note that we project $\bm{\eta}^{+(t)}$ and $\bm{\eta}^{-(t)}$ such that $\eta_{1}^{+(t)}=\eta_{1}^{-(t)}=0$ when $K=2$, which are still denoted as $\bm{\eta}^{+(t)}$ and $\bm{\eta}^{-(t)}$ with a slight abuse of notation.
Next, we adopt the similar method as in yuan2021community and zhang2024consistent to update $\bm{Z}^{(t)}$ due to its discrete nature. Let $\bm{Z}_{-i}$ be a matrix by deleting of the ii‑th row of $\bm{Z}$. Then, node $i$ will be more likely to belong to the kk‑th community if $L(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z}_{-i},z_{ik}=1)\leq L(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z}_{-i},z_{il}=1)$ for each $l\neq k$. Thus, we have
$z_{ik}^{(t)}=\begin{cases}1,&k=\displaystyle\arg\min_{l\in[K]}L\left(\bm{\gamma}^{+(t)},\bm{\eta}^{+(t)},\bm{\gamma}^{-(t)},\bm{\eta}^{-(t)},\bm{Z}_{-i}^{(t-1)},z_{il}=1\right);\\ 0,&\text{otherwise},\end{cases}$
for each $i\in[n]$ and $k\in[K]$.
After updating $\bm{Z}^{(t)}$, we check whether
$L\left(\bm{\gamma}^{+(t)},\bm{\eta}^{+(t)},\bm{\gamma}^{-(t)},\bm{\eta}^{-(t)},\bm{Z}^{(t)}\right)<L\left(\bm{\gamma}^{+(t)},\bm{\eta}^{+(t)},\bm{\gamma}^{-(t)},\bm{\eta}^{-(t)},\bm{Z}^{(t-1)}\right).$
Otherwise, we update each $\bm{z_{i}}$ sequentially, thereby ensuring that the negative log‑likelihood always decreases. Let $\alpha$ and $T_{\max}$ be threshold values for the relative error and updating steps. The whole algorithm will stop when $t>T_{\max}$ or the relative error
$\frac{\bigl|L(\bm{\gamma}^{+(t)},\bm{\eta}^{+(t)},\bm{\gamma}^{-(t)},\bm{\eta}^{-(t)},\bm{Z}^{(t)})-L(\bm{\gamma}^{+(t-1)},\bm{\eta}^{+(t-1)},\bm{\gamma}^{-(t-1)},\bm{\eta}^{-(t-1)},\bm{Z}^{(t-1)})\bigr|}{\bigl|L(\bm{\gamma}^{+(t-1)},\bm{\eta}^{+(t-1)},\bm{\gamma}^{-(t-1)},\bm{\eta}^{-(t-1)},\bm{Z}^{(t-1)})\bigr|}<\alpha.$
Let $\widehat{\bm{\gamma}}^{+}$, $\widehat{\bm{\eta}}^{+}$, $\widehat{\bm{\gamma}}^{-}$, $\widehat{\bm{\eta}}^{-}$ and $\widehat{\bm{Z}}$ be the output of the alternative updating scheme, then $\widehat{\bm{\Theta}}^{+}=\widehat{\bm{\gamma}}^{+}\bm{1}_{n}^{\top}+\bm{1}_{n}\widehat{\bm{\gamma}}^{+\top}+(\widehat{\bm{\eta}}^{+}\bm{1}_{n}^{\top}+\bm{1}_{n}\widehat{\bm{\eta}}^{+\top})\odot\widehat{\bm{Z}}\widehat{\bm{Z}}^{\top}$ and $\widehat{\bm{\Theta}}^{-}=\widehat{\bm{\gamma}}^{-}\bm{1}_{n}^{\top}+\bm{1}_{n}\widehat{\bm{\gamma}}^{-\top}+(\widehat{\bm{\eta}}^{-}\bm{1}_{n}^{\top}+\bm{1}_{n}\widehat{\bm{\eta}}^{-\top})\odot\widehat{\bm{Z}}\widehat{\bm{Z}}^{\top}$.

### v2

@@ Abstract
Network data has attracted growing interest across scientific domains, prompting the development of various network models. Existing network analysis methods mainly focus on unsigned networks, whereas signed networks, consisting of both positive and negative edges, have been frequently encountered in practice but much less investigated. In this paper, we formally define strong and weak balance in signed networks, and propose a signed block $\beta$-model, which is capable of modeling strong- and weak-balanced signed networks simultaneously. We establish the identifiability of the proposed model by leveraging properties of bipartite graphs, and develop an efficient two-step algorithm to tackle the resulting optimization problem and recover the community membership. More importantly, we establish the asymptotic consistencies of the proposed model in terms of both parameter estimation and community detection. Its advantages are also demonstrated through extensive numerical experiments and the application to a real-world international relationship network.
@@ 1 Introduction
Network data has attracted increasing attention in recent years, with applications in diverse scientific domains, ranging from social science [Hunter et al., 2008], computer science [Heard et al., 2014] to biomedicine [Shojaie and Michailidis, 2009]. Analysis of network data can reveal latent structures underlying the interactions among entities of interest, such as community detection [Girvan and Newman, 2002], link prediction [Martínez et al., 2016], homophily [Hoff, 2007] and so on.
Most existing methods for network analysis focus on unsigned networks, whereas signed networks are also prevalent in real-world applications. Specifically, signed networks have been extensively employed to represent relationships among objects [Harary, 1953], which associate polarity information with each edge and thus distinguish friends from enemies, or allies from rivals. These signed edges give rise to unique structural characteristics grounded in social psychology theories, particularly the balance theory [Heider, 1946; Cartwright and Harary, 1956], stating that certain triadic configurations involving signed edges lead to stable or unstable social structures. Figure 1 illustrates four possible triads, among which triads A and C represent strong-balanced structures. These triads align with intuitive social principles, such as “a friend of my friend is likely my friend” and “an enemy of my friend is likely my enemy”. Weak balance [Davis, 1967] extends this concept by additionally permitting triad D, which accommodates the occurrence of three mutually hostile nodes and is thus more suitable for modeling real-life signed networks.
@@
Community detection has been widely studied for unsigned networks [Holland et al., 1983; Girvan and Newman, 2002; Newman, 2006; Lei and Rinaldo, 2015; Zhang et al., 2022a], but much less so for signed networks. Most existing community detection methods for signed networks [Kunegis et al., 2010; Jiang, 2015; Cucuringu et al., 2021] are algorithm based and lack formal theoretical analysis, not to mention that the prevalent balanced structures in signed networks are often ignored by the aforementioned community detection methods. Note that the balance theory actually posits that a strong-balanced (weak-balanced) network can be partitioned into two (multiple) sub-networks such that positive edges connect nodes within the same sub-network, whereas nodes from different sub-networks are connected by negative edges [Easley et al., 2010]. More recently, tang2024population introduces a formal definition of strong-balanced signed networks in a probabilistic sense and proposes an efficient method to detect the two communities, yet the method is inapplicable to analyze the more realistic weak-balanced signed networks with multiple communities.
@@
In this paper, we first provide a new set of conditions for both strong- and weak-balanced signed networks, from both local and global perspectives. Built upon the global perspective, we then propose a signed block $\beta$-model to explicitly incorporate the strong or weak balance structures in heterogeneous signed networks. Particularly, it introduces two sets of parameters to capture each node’s distinct connection tendencies within its own community and across different communities. The identifiability of the signed block $\beta$-model is established by exploiting properties of the bipartite graph, and an efficient two-step algorithm is developed to solve the resulting optimization problem and recover the community membership. More importantly, the asymptotic consistencies of the signed block $\beta$-model are established in terms of both parameter estimation and community detection. Its advantages are also supported in numerical experiments on various synthetic signed networks and a real-world international relationship network.
@@
The main contribution of the proposed signed block $\beta$-model is that it provides a novel framework to statistically model strong and weak balance structures in signed networks. It also generalizes the traditional $\beta$-model [chatterjee2011random; 10.1214/12-AOS1078; chen2021analysis] by incorporating community structures while preserving its capacity to capture node heterogeneity. For each node, two sets of parameters are introduced to model its tendency to positively or negatively connect with other nodes within the same community or across different communities. Such a treatment not only incorporates community structure in each node’s connection tendency, but also admits more substantial node heterogeneity than existing models [chen2021analysis; shao20212; cucuringu2021regularized]. In addition, a novel set of identifiability conditions for the proposed signed block $\beta$-model is established by leveraging techniques from the graph theory. The theoretical techniques developed here can be applicable to identifiability analyses of other network models involving community structures, and thus are of independent interest. Moreover, we incorporate the $(1+\epsilon)$-approximation algorithm from the literature of projective clustering har2002projective; deshpande2006matrix; tukan2022new into the second step of our method to solve the $K$-Lines clustering problem, namely, to compute $K$ lines to approximate a set of points. Although li2024dissecting discusses the $K$-Lines clustering problem, their algorithm lacks theoretical guarantee.
@@
The remainder of this paper is organized as follows. Section 2 formally gives the definition of strong and weak balance in signed networks. Section 3 presents the signed block $\beta$-model, establishes its identifiability conditions, and develops an efficient two-step algorithm to tackle the resulting optimization task. Section 4 provides a rigorous theoretical analysis of the proposed signed block $\beta$-model and quantifies its asymptotic behaviors in both parameter estimation and community detection. Section 5 examines the numeric performance of the signed block $\beta$-model via extensive experiments on both synthetic and real-world signed networks. Finally, Section 6 concludes the paper with a brief summary. Technical proofs of the main theorems and propositions are provided in the Appendix, whereas auxiliary lemmas are contained in the Supplementary Material.
@@
We introduce some notations to be used throughout the paper. For any integer $n$, let $n]=\{1,\dots,n\}$. For any vector $\bm{v}=(v_{1},\dots,v_{n})^{\top}\in\mathcal{R}^{n}$, $\|\bm{v}\|_{2}^{2}=\sum_{i\in[n]}v_{i}^{2}$, and $\|\bm{v}\|_{\infty}=\max_{i\in[n]}|v_{i}|$. We use $\bm{1}_{n}$ to denote the $n$-dimensional column vector with all elements being 1 and $\mathbf{1}(\cdot)$ to denote the indicator function. For any matrix $\bm{A}=(a_{ij})\in\mathcal{R}^{n\times m}$, let $\bm{A}_{i\cdot}$ be its $i$-th row, $\bm{A}_{\cdot i}$ be its $i$-th column, and $\|\bm{A}\|_{F}^{2}=\sum_{i\in[n],j\in[m]}a_{ij}^{2}$ be its Frobenius norm. For any two matrices $\bm{A}$ and $\bm{B}$, $\bm{A}\odot\bm{B}$ is the Hadamard product with $(\bm{A}\odot\bm{B})_{ij}=a_{ij}b_{ij}$. Let $\bm{J}=I_{n}-\frac{\bm{1}_{n}\bm{1}_{n}^{\top}}{n}$ be the centering matrix. For any two vectors $\bm{x}=(x_{1},\dots,x_{n})$ and $\bm{y}=(y_{1},\dots,y_{n})$, $\bm{x}<\bm{y}$ denotes $x_{i}<y_{i}$ for each $i\in[n]$. For any two sequences $\{a_{n}\}_{n\geq 1}$ and $\{b_{n}\}_{n\geq 1}$, we denote $a_{n}\prec b_{n}$ if $a_{n}=o(b_{n})$, $a_{n}\preceq b_{n}$ if $a_{n}=O(b_{n})$ and $a_{n}\simeq b_{n}$ if $a_{n}\preceq b_{n}$ and $b_{n}\preceq a_{n}$. For any set $\mathcal{X}$, $|\mathcal{X}|$ denotes its cardinality and $\mathcal{X}^{c}$ denotes its complement. For any two sets $\mathcal{X}$ and $\mathcal{Y}$, $\mathcal{X}\setminus\mathcal{Y}=\{x\mid x\in\mathcal{X}\text{ and }x\notin\mathcal{Y}\}$.
2 Signed network and balance structure
@@
Definition 1 (Strong balance).
@@
Definition 1 assures that a strong-balanced signed network does not contain a triad achieving the highest probability at triad B or D as shown in Figure 1. It can also be extended to define weak-balanced signed networks.
Definition 2 (Weak balance).
@@
Propositions 1 and 2 characterize the strong and weak balance structures in signed networks from both local and global perspectives. In fact, the local perspective (i) in Proposition 1 is essentially equivalent to $\mathbb{E}(a_{ij}a_{jk}a_{ki}\big||a_{ij}a_{jk}a_{ki}|=1)>0$ for any $(i,j,k)$, which is exactly the definition of the strong-balanced signed network in tang2024population, whereas Proposition 2(i) can be regarded as its analogy for the weak-balanced signed network. On the other hand, Proposition 1(ii) and 2(ii) give the brand new global perspectives on strong and weak balance structures based on partitions of the signed network, which greatly facilitate the modeling of community structures in signed networks.
@@ 3 Signed block $\beta$-model
We propose a signed block $\beta$-model, denoted as SBBM, to leverage the global perspectives of the balance structures in signed networks. Particularly, let $\sigma:[n]\rightarrow[K]$ denote the community membership of node $i$, then
@@
It can be readily verified that nodes $i$ and $j$ belong to the same community if and only if $\mathbb{P}(a_{ij}=1)>\mathbb{P}(a_{ij}=-1)$, and different communities if and only if $\mathbb{P}(a_{ij}=1)<\mathbb{P}(a_{ij}=-1)$. It follows from Propositions 1 and 2 that SBBM with constraints (2) satisfies strong balance when $K=2$ and weak balance when $K\geq 3$. In contrast, enforcement of strong and weak balance structures can be challenging in most other existing signed network models.
@@ 3.1 Model identifiability
Let $\bm{Z}=(\bm{z}_{1},\dots,\bm{z}_{n})^{\top}$ be the community membership matrix, where $\bm{z}_{i}\in\{0,1\}^{K}$ with $z_{ik}=1$ if $\sigma(i)=k$. Define $\bm{\beta}^{+}=(\beta^{+}_{i})_{i=1}^{n}$, $\bm{\gamma}^{+}=(\gamma_{i}^{+})_{i=1}^{n}$, $\bm{\beta}^{-}=(\beta^{-}_{i})_{i=1}^{n}$, $\bm{\gamma^{-}}=(\gamma_{i}^{-})_{i=1}^{n}$, $\bm{\eta}^{+}=\bm{\beta}^{+}-\bm{\gamma}^{+}$ and $\bm{\eta}^{-}=\bm{\beta}^{-}-\bm{\gamma}^{-}$. Then, the parameter space is
$\mathcal{P}=\big\{(\bm{\gamma}^{+},\bm{\eta}^{+},\bm{\gamma}^{-},\bm{\eta}^{-},\bm{Z})\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}$ $\times\mathbb{R}^{n}\times\{0,1\}^{n\times K}:$
@@ 3.2 Computing algorithm
The negative log-likelihood of $\bm{A}^{+}=(a^{+}_{ij})_{i,j=1}^{n}$ and $\bm{A}^{-}=(a^{-}_{ij})_{i,j=1}^{n}$ can be written as
$\mathcal{L}_{n}(\bm{\Theta}^{+},\bm{\Theta}^{-})=-\sum_{1\leq i,j\leq n}\Big(a^{+}_{ij}\theta^{+}_{ij}+a^{-}_{ij}\theta^{-}_{ij}-\log\big(1+e^{\theta^{+}_{ij}}+e^{\theta^{-}_{ij}}\big)\Big),$
where $a_{ij}^{+}=\mathbb{I}(a_{ij}=1)$ and $a_{ij}^{-}=\mathbb{I}(a_{ij}=-1)$. Then we have
$(\widehat{\bm{\Theta}}^{+},\widehat{\bm{\Theta}}^{-})=\operatorname*{arg\,min}_{\bm{\Theta}^{+},\bm{\Theta}^{-}}\left\{\mathcal{L}_{n}(\bm{\Theta}^{+},\bm{\Theta}^{-})+\lambda_{n}^{+}\|\bm{\Theta}^{+}\|_{*}+\lambda_{n}^{-}\|\bm{\Theta}^{-}\|_{*}\right\},$ (3)
where $\|\cdot\|_{*}$ denotes the nuclear norm, which is used to promote the low-rank structure in $\bm{\Theta}^{+}$ and $\bm{\Theta}^{-}$, and $\lambda_{n}^{+}$ and $\lambda_{n}^{-}$ are two tuning parameters. The objective function in (3) is convex, and it can be efficiently optimized via the accelerated proximal gradient algorithm [beck2009fast]. The algorithm iterates between a gradient descent step on the smooth loss $\mathcal{L}_{n}$ and a proximal step for the nuclear norm, which is computed via the singular value thresholding operator [cai2010singular], $\mathcal{D}_{\tau}(\bm{M})=\sum_{k}\max\{\sigma_{k}-\tau,0\}\bm{u}_{k}\bm{v}_{k}^{\top}$ for any matrix $\bm{M}$ with singular value decomposition $\bm{M}=\sum_{k}\sigma_{k}\bm{u}_{k}\bm{v}_{k}^{\top}$. Once $\widehat{\bm{\Theta}}^{+}$ and $\widehat{\bm{\Theta}}^{-}$ are obtained, we then proceed to estimate $\sigma$.
Proposition 3.
Let $\bm{\Theta}_{\Delta}=\bm{\Theta}^{+}-\bm{\Theta}^{-}$ and $\bm{J}\bm{\Theta}_{\Delta}\bm{J}=\bm{Y}\bm{\Sigma}\bm{Y}^{\top}$ be the eigen decomposition with $\bm{Y}\in\mathbb{R}^{n\times r}$ and $r=\text{rank}\ (\bm{J}\bm{\Theta}_{\Delta}\bm{J})$. Then there are $K$ lines in $\mathbb{R}^{r}$, denoted as $\{l_{k}:\bm{w}_{k}+t\bm{v}_{k},t\in\mathbb{R},\bm{w}_{k},\bm{v}_{k}\in\mathbb{R}^{r},k=1,\dots,K\}$, such that each row of $\bm{Y}$ lies exactly on one of these $K$ lines. Furthermore, for any $i,j\in[n]$, if $\sigma(i)=\sigma(j)\in[K]$, then $\bm{Y}_{i\cdot}$ and $\bm{Y}_{j\cdot}$ live on the same line.
Proposition 3 implies that the community structure in $\sigma$ can be estimated from the spectral decomposition of $\bm{J}\widehat{\bm{\Theta}}_{\Delta}\bm{J}$, where $\widehat{\bm{\Theta}}_{\Delta}=\widehat{\bm{\Theta}}^{+}-\widehat{\bm{\Theta}}^{-}$. Denote the eigenvectors of $\bm{J}\widehat{\bm{\Theta}}_{\Delta}\bm{J}$ as $\widehat{\bm{Y}}=(\widehat{\bm{Y}}_{1},\ldots,\widehat{\bm{Y}}_{n})^{\top}$, then we can recover $\sigma$ via a $K$-lines clustering of $\widehat{\bm{Y}}$. Formally, given $\{\widehat{\bm{Y}}_{1},\ldots,\widehat{\bm{Y}}_{n}\}$ and a set of lines $\bm{L}=\{\ell_{1},\dots,\ell_{K}\}$, we define
$\Phi(\widehat{\bm{Y}},\bm{L},\sigma)=\frac{1}{n}\sum_{i=1}^{n}\|\widehat{\bm{Y}}_{i}-P_{\ell_{\sigma(i)}}\widehat{\bm{Y}}_{i}\|_{2}^{2},$
where $P_{l}$ indicates the projection onto the $l$-th line. Let $\sigma_{\bm{L}}(i)=\arg\min_{k\in[K]}\|\widehat{\bm{Y}}_{i}-P_{l_{k}}\widehat{\bm{Y}}_{i}\|_{2}$, then $\Phi(\bm{Y},\bm{L},\sigma_{\bm{L}})=\min_{\sigma}\Phi(\bm{Y},\bm{L},\sigma)$. Therefore, the $K$-line clustering aims to find $\widehat{\bm{L}}=\arg\min_{\bm{L}}\Phi(\widehat{\bm{Y}},\bm{L},\sigma_{\bm{L}})$, which becomes a projective clustering problem [tukan2022new]. The general projective clustering problem is NP-hard, but various techniques have been developed to obtain its $(1+\epsilon)$-approximated solutions [feldman2006coresets; ding2012linear; feldman2020turning]. In particular, we adopt the method in ding2012linear, which guarantees an objective value no greater than $(1+\epsilon)$ times the optimal value over the entire search space. Denote the solution as $\widehat{\bm{L}}=\{\widehat{\ell}_{1},\dots,\widehat{\ell}_{K}\}$ such that
$\Phi(\widehat{\bm{Y}},\widehat{\bm{L}},\sigma_{\widehat{\bm{L}}})\leq(1+\epsilon)\,\min_{\bm{L}}\Phi(\widehat{\bm{Y}},\bm{L},\sigma_{\bm{L}}),$
and then the community membership is estimated as $\widehat{\sigma}=\sigma_{\widehat{\bm{L}}}$.